# A Study of Undefined Behavior Across Foreign Function Boundaries in Rust Libraries

### v4

@@ Abstract
Developers rely on the Rust programming language’s static safety guarantees to write secure and performant applications. However, Rust is frequently used to interoperate with other languages which allow design patterns that conflict with Rust’s aliasing models. Miri is the only dynamic analysis tool capable of validating applications against these models, but it does not support foreign functions, indicating that there may be a critical correctness gap at the heart of the Rust ecosystem. We conducted a large-scale evaluation of Rust libraries that call foreign functions to determine whether Miri’s dynamic analyses remain useful in this context. We used Miri and an LLVM interpreter to jointly execute applications that call foreign functions, where we found 48 instances of undefined or undesired behavior. These include three bugs from libraries that had over 10,000 daily downloads on average during our observation period and one from a library maintained by the Rust Project. Many of the errors we found involved incompatible aliasing patterns, but Rust’s latest Tree Borrows aliasing model was significantly more permissive than the earlier Stacked Borrows model. The Rust community must invest in new, production-ready tooling for multi-language applications to ensure that developers can detect these errors.
I Introduction
The Rust programming language has been increasingly popular due to its static safety guarantees, which provide security benefits comparable to garbage-collection without additional run-time overhead [1, 2]. However, when Rust’s aliasing restrictions become burdensome, developers can bypass them using a subset of unsafe features. If these features are used incorrectly, they can break the rules of Rust’s aliasing model. The Rust compiler relies on these rules to optimize code. If they are broken, optimizations may be applied incorrectly, which can introduce security vulnerabilities.
Miri [3] is a widely-used Rust interpreter that detects violations of the language’s aliasing model via dynamic analysis. A limitation of Miri is that it cannot detect these violations across foreign function boundaries. However, one of the most common reasons for using unsafe features is to call foreign functions [4, 5, 6, 7, 8]. We seek to determine whether Miri’s dynamic checking would still be useful in this context.
RQ1: Are Miri’s methods for dynamically enforcing Rust’s aliasing model useful for detecting undefined behavior in programs that call foreign functions?
@@ I Introduction
[…]
@@
In Section II, we compare Rust’s semantics with C and C++, and we describe the resources and best-practices that Rust developers can use to ensure correct interoperation. In Section III, we document our methodology for sampling and evaluating test cases from Rust libraries that call foreign functions, and we describe the challenges that we encountered when implementing MiriLLI. In Section IV, we describe each type of bug that we found. We discuss the implications of our findings in Section V. We review prior work on Rust interoperation in Section VI, we discuss threats to validity in Section VII, and we conclude in Section VIII. Our dataset, Appendix, and the source for MiriLLI are available through GitHub111https://github.com/icmccorm/mirilli.
@@ II Background
Rust’s safety restrictions begin at the level of a value, which is valid for a particular scope. The type of a value implements “traits” which define its behavior. For example, all types have “move semantics” by default, meaning that each value has a unique “owner.” However, values with the Copy trait have no owner and can be freely duplicated. Ownership can be transferred through assignment or borrowed by creating a reference. References are either mutable, taking the form &mut T, or immutable, with the form &T. Mutable references have move semantics, but immutable references can be copied. Each reference has a “lifetime,” which is the portion of the program over which it will be used. Rust’s borrow checker statically enforces that the lifetime of a reference must not exceed the scope or lifetime of the value that it refers to. A value can only have a single mutable reference or many immutable references that are active within a given context [12], but not both at the same time.
These restrictions eliminate safety issues, but they can be overly conservative for certain design patterns. When developers need to bypass Rust’s restrictions, they can use the unsafe keyword to enable a set of additional features that are not restricted by the borrow checker. These features include dereferencing raw pointers, accessing the fields of union types, modifying static mutable state, implementing unsafe traits, and calling unsafe functions—including those written in other languages [13]. The Rust compiler assumes that programs using unsafe code will follow the rules of its aliasing model. Programs that break these rules have undefined behavior and may be optimized incorrectly. This can lead to differences in the behavior that have introduced security vulnerabilities [14] in practice.
@@ II-A Rust’s Aliasing Model
Rust developers can avoid these issues by ensuring that their programs adhere to the rules of Rust’s aliasing model. The Tree Borrows [10] model provides the latest definition of these rules. Under this model, a pointer is both an address to a location in memory and a permission to read or write to that location. Each location is associated with a tree that tracks all valid permissions to its contents. When a location is borrowed, a new branch is created in the tree carrying a new permission to access the location. When a pointer is used to access a location, its permission is compared with the current state of the tree to ensure that it is valid and that it permits the kind of access taking place. Each type of access will have a different effect on the tree depending on the position of each permission in the tree relative to the permission used for the access. With respect to a given permission, an access is a considered a child access if requires that permission or any of its descendants; otherwise, it is a foreign access.
It is undefined behavior to access a location using an invalid permission. Figure 1 illustrates the three categories of undefined behavior under Tree Borrows. The first category is an expired permission error, which is shown in Figure 1(a). When the variable x is declared on line 1, a new tree is created containing a permission, labelled . The permission allows read and write accesses, and it is the default state for a mutable variable. On line 2, a mutable borrow of x is assigned to y, which creates an additional, adjacent branch in the tree with the permission . This type of permission behaves like a read-only permission, but it becomes when it is used for a write access.
On line 3, the reference assigned to y is cast into a raw pointer, which still carries permission . Then, the borrow assigned to z on line 4 creates another branch in the tree with permission . Next, y is mutated line 5, causing permission to transition from to . This is a foreign access relative to permission , so its permission transitions to . The permission does not permit any kind of access; this transition enforces the uniqueness of mutable references. Since permission is now , the write access through z on line 6 is undefined behavior. We refer to this as an expired permission error.
@@
The second type of undefined behavior is an insufficient permission error, which occurs when a valid permission exists in the tree but does not permit the type of access taking place. This occurs in Figure 1(b), where the function mutate writes through a mut pointer cast from an immutable reference. Though the cast is valid, the pointer still has the read-only, permission created by the immutable reference, so writing through it is undefined behavior.
After finding a valid, “granting” permission [9], the tree is modified based on the type of access taking place. A write access may cause permissions to become , as shown in Figure 1(a). However, the permissions associated with reference-type arguments become “protected” upon entering the body of a function. It is undefined behavior for a protected tag to become before its associated call returns. When a function returns, the permissions of its arguments are no longer protected. We demonstrate this in Figure 1(c), where deallocating through x on line 2 causes its permission to become . Since x is a reference-type argument, it is protected, so this is undefined behavior. We refer to an access that disables a protected tag as a framing error.
@@ II-A1 Differences from Stacked Borrows
Tree Borrows replaces the Stacked Borrows [9] model, which uses a stack to track the active permissions to each location. Under Stacked borrows, a reference cannot be offset and used to access locations outside of the range it originally borrowed from. Figure 2 shows an example of this invalid range error; the dereference on line 2 is invalid since x only borrows the first element of the tuple. This pattern is allowed under Tree Borrows; a valid permission can be used to access any location within bounds. Stacked Borrows may also eagerly invalidate permissions when a new reference is created, even if it is never used. Under Tree Borrows, only memory accesses will trigger undefined behavior. However, Tree Borrows’ handling of mutable references is more strict than Stacked Borrows in situations involving “two-phase” borrowing, where a unique mutable reference is treated immutable until its first write access. Tree Borrows models this consistently with the permission, while Stacked Borrows only applies these semantics in certain situations.
@@
Rust shares several other categories of undefined behavior with current C [15] and C++ [16] standards, such as accessing memory that has been freed, accessing beyond the bounds of an allocation, creating a data race, reading uninitialized memory, and accessing a value at an unaligned address. However, C and C++ have different rules related to initialization, aliasing, and mutability that may conflict with Rust’s expectations.
@@ II-B1 Initialization
In both C and C++, accessing an object in memory before it has been initialized produces an “indeterminate value”, which is undefined behavior222C23 §6.2.4.6-7; C++23 §6.7.4.1. There are a few exceptions, however; static and thread-local variables are initialized immediately at the beginning of their lifetime, and both unsigned characters and bytes can soundly store indeterminate values. It is also safe to create a reference to a value that is uninitialized. The initialization of a value is not encoded into the type system by default; developers must manually reason about the state of memory or rely on instrumentation to detect uninitialized reads.
Rust’s borrow checker prevents variables from being used until they are initialized, and no primitive types tolerate indeterminate values in safe contexts. However, Rust’s standard library provides the MaybeUninit<T> struct, which represents an instance of T that may or may not be initialized. It is undefined behavior to produce a safe reference to uninitialized memory, so the macros addr_of! and addr_of_mut! must be used to obtain raw pointers to the interior of MaybeUninit<T> without borrowing it. Once an instance has been fully initialized, the unsafe function MaybeUninit<T>::assume_init() will unwrap the outer struct to produce the value inside. It is undefined behavior to call assume_init() if T is not fully initialized.
@@ II-B2 Aliasing & Mutability
Neither C nor C++ statically restrict aliasing, but the standards for each language allow implementations to make type-directed assumptions about aliasing. In both languages, it is considered undefined behavior for a pointer to one type to refer to a value of another type if the two types differ333C23 §6.5.7; C++23 §7.2.1.11. However, differences do not include qualifiers, so two parameters of types int * and const int * may alias. Developers writing C programs can use the restrict type qualifier to hint to the compiler that two pointers with compatible types should never alias444C23 §6.7.3.9, but compilers are not obliged to consider restrict when performing optimizations.
In Rust, both variables and references have distinct, connected capabilities for mutation. Variables are immutable by default, only mutable variables can be mutably borrowed, and only mutable references can be used for mutation. In C and C++, variables are mutable by default, and the capability to mutate memory is determined by the mutability of the object being pointed to, regardless of the pointer’s type555C23 §6.7.3.7; C++23 §7.6.1.11.6. If a variable was declared as mutable, then a const pointer can be cast into a mutable pointer and used to write to it.
@@ II-B3 Provenance
Both C and C++ allow pointers to be converted to and from integers with a size equal to the word size of the current architecture666C23 §6.3.2.3.5-6; C++23 §6.7.1.10.4-5. Implementations may track the “provenance,” or origin of pointers to inform compile-time optimizations [17], but neither of the current standards for C or C++ define if or how provenance should be preserved across these conversions. The Rust project has settled on the notion of pointers having provenance [18], but the specifics of Rust’s provenance model are still being decided.
Miri has multiple methods of handling provenance across pointer-to-integer conversion. By default, Miri uses a “permissive provenance” model, which matches Memarian et al. [19]’s “PVNI-wildcard” model. When a pointer is converted into an integer, its tag is added to a set of tags that have been “exposed” by pointer-to-integer conversion. When integers are converted back into pointers, their allocation identifier is reconstructed from their address, but they receive a “wildcard” provenance value. Under Stacked Borrows, when a memory access occurs using a wildcard tag, Miri’s eagerly interprets it as being equivalent to any tag in the stack that permits the access. Tree Borrows does support this behavior yet; all accesses are allowed through wildcard tags without modifying the tree. Alternatively, developers can enable “strict provenance”, which treats integer to pointer conversion as an error.
@@ II-C Interoperation
Developers can access foreign functions and static variables using Rust’s foreign function interface. Declarations, or “bindings”, for each of these objects are written in extern blocks, which take an optional qualifier to indicate the Application Binary Interface (ABI) used by the foreign library. Only a subset of Rust’s types are guaranteed to be compatible with foreign ABIs, and both structs and enums must be annotated with the #[repr(C)] attribute to ensure that their layout is compatible with C. The Rust compiler enforces the use of ABI-compatible types via lints, but these can be disabled. Bindings are not validated against their definitions, but tools such as Bindgen [20] and CXX [21] can automatically generate bindings from existing codebases.
Several dynamic analysis tools can find undefined behavior in multi-language Rust applications. Valgrind [22] is capable of flagging spatial and temporal memory errors, as well as incompatibilities with size and alignment. Many of the LLVM Project’s sanitizers are also compatible with Rust [23]. Each of these tools is available as a plugin for Cargo, Rust’s build tool. Miri can call foreign functions from natively compiled shared libraries using libffi [24], but only integer and floating-point arguments are supported. None of these tools can validate foreign libraries against Rust’s aliasing model.
@@ III Methodology
We seek to determine whether the differences between Rust and other languages lead to undefined behavior in practice. First, we analyzed all libraries published on Rust’s central package repository to find the subset with test cases that produce LLVM bitcode for C or C++ libraries during their build process; we describe this stage in Section III-A. Then, we created MiriLLI to analyze runs of these test cases for undefined behavior. MiriLLI extends Miri to interoperate with LLI [25], an LLVM interpreter, so that both interpreters can jointly execute programs defined across LLVM’s intermediate representation (LLVM IR) and Rust’s Middle Intermediate Representation (MIR). We describe the architecture of our tool in Section III-B, and we account for several changes that were necessary to resolve differences in semantics between each interpreter. In Section III-C, we describe our method for deduplicating test results to identify unique instances of undefined behavior and our approach to reporting bugs.

### v5

@@ Abstract
Developers rely on the static safety guarantees of the Rust programming language to write secure and performant applications. However, Rust is frequently used to interoperate with other languages which allow design patterns that conflict with Rust’s evolving aliasing models. Miri is currently the only dynamic analysis tool that can validate applications against these models, but it does not support foreign functions, indicating that there may be a critical correctness gap across the Rust ecosystem. We conducted a large-scale evaluation of Rust libraries that call foreign functions to determine whether Miri’s dynamic analyses remain useful in this context. We used Miri and an LLVM interpreter to jointly execute applications that call foreign functions, where we found 47 instances of undefined or undesired behavior from 37 libraries. Three bugs were found in libraries that had more than 10,000 daily downloads on average during our observation period, and one was found in a library maintained by the Rust Project. Many of these bugs were violations of Rust’s aliasing models, but the latest Tree Borrows model was significantly more permissive than the earlier Stacked Borrows model. The Rust community must invest in new, production-ready tooling for multi-language applications to ensure that developers can detect these errors.
Index Terms:
Rust, C, C++, interoperation, undefined behavior, aliasing, bugs, foreign functionsI Introduction
The Rust programming language has become increasingly popular due to its static safety guarantees, which provide security benefits comparable to garbage collection without additional run-time overhead [1, 2]. However, Rust is also frequently used in interoperation with languages that do not provide similar assurances. To call foreign functions, developers must use a subset of unsafe features to bypass Rust’s restrictions. If these features are used incorrectly, they can break Rust’s aliasing rules. The Rust compiler relies on these rules to optimize code. If they are broken, optimizations may be applied incorrectly, which can introduce security vulnerabilities.
Miri [3] is a widely-used Rust interpreter that uses dynamic analysis to detect violations of Rust’s aliasing model. A limitation of Miri is that it cannot detect these errors across foreign function boundaries. However, one of the most common reasons for using Rust’s unsafe features is to call foreign functions [4, 5, 6, 7, 8]. We seek to determine whether the differences between Rust and other languages can lead to errors in practice.
RQ1: What types of errors occur in Rust libraries that call foreign functions?
@@
To answer these questions, we created MiriLLI: a tool which combines Miri with an LLVM interpreter to jointly execute programs and detect undefined behavior across foreign function boundaries. We used MiriLLI to conduct a large-scale study of 9,130 test cases from 957 Rust libraries that call foreign functions. We identified 47 instances of undefined or undesirable behavior in 37 of these libraries. Of the 90 test cases that violated Stacked Borrows, 66% (59) did not violate Tree Borrows.
@@
In Section II, we compare Rust’s semantics with C and C++, and we describe the resources and best practices that Rust developers use to interoperate with these languages. In Section III, we document our methodology for sampling and evaluating test cases from Rust libraries that call foreign functions, and we describe the challenges that we encountered when implementing MiriLLI. In Section IV, we describe each type of bug that we found. We discuss the implications of our findings in Section V. We review prior work on Rust interoperation in Section VI, we discuss threats to validity in Section VII, and we conclude in Section VIII. Our dataset, the Appendix, and the source for MiriLLI are available in our replication package111https://github.com/icmccorm/mirilli.
@@ II Background
Rust’s safety restrictions begin at the level of a value, which is valid for a particular scope. A value’s type implements traits that define how it behaves. For example, all types have move semantics by default, meaning that each value has a unique owner. However, values with the Copy trait have no owner and can be freely duplicated. Ownership can be transferred through assignment or borrowed by creating a reference. References are either mutable, taking the form &mut T, or immutable, with the form &T. Mutable references have move semantics, but immutable references can be copied. Each reference has a lifetime, which is the portion of the program over which it will be used. Rust’s borrow checker statically enforces that the lifetime of a reference must not exceed the scope or lifetime of the value that it borrows. A value can have a single mutable reference or many immutable references that are active within a given context [12], but not both at the same time.
These restrictions prevent safety issues, but they can also make it impossible to implement certain design patterns. When developers need to bypass Rust’s restrictions, they can use the unsafe keyword to enable a set of additional features that are not restricted by the borrow checker. These features include dereferencing raw pointers, accessing the fields of union types, modifying static mutable state, implementing unsafe traits, and calling unsafe functions—including those written in other languages [13]. The Rust compiler can assume that programs that use unsafe code will follow the rules of its aliasing model. Programs that violate these rules have undefined behavior, so they may be optimized incorrectly. This can lead to differences in behavior that can cause security vulnerabilities [14] in practice.
@@ II-A Rust’s Aliasing Model
Rust developers can avoid these issues by ensuring that their programs adhere to the rules of Rust’s aliasing model. The Tree Borrows [10] model provides the latest definition of these rules. Under this model, each pointer has provenance that determines its permission to read or write to a location [15]. Each location is associated with a tree that tracks all valid permissions to its contents. When a location is borrowed, a branch is created in the tree that holds a new permission to access the location. When a pointer is used, its permission is checked against the tree to ensure that it is valid and that it permits the kind of access that takes place. Each type of access will have a different effect on the tree depending on the position of each permission in the tree relative to the permission used for the access. With respect to a given permission, an access is considered a child access if it requires that permission or any of its descendants; otherwise, it is a foreign access. Accessing a location using an invalid permission is considered undefined behavior.
We created three categories to classify the Tree Borrows violations that we observed in our investigation. Figure II-A provides minimal examples of each one. We refer to the first category as an expired permission error. Under Tree Borrows, writing to a memory location may cause other permissions to that location to expire (or transition to , in terms of the model). Using one of these expired permissions is undefined behavior. We provide a minimal example of this in Figure 2, where the variable x is mutably borrowed twice; once by y and again by z. However, we also store a copy of y as a raw pointer (*mut). Rust’s borrow checker does not place aliasing restrictions on raw pointers, so it assumes that z has exclusive, mutable access from when it is created on line 4 until it is used on line 6. Writing through y on line 5 breaks this assumption, so writing through z is now undefined behavior. In terms of the model, the permissions for y and z are children of the permission for x. Both receive a permission, which has the capability to become a unique permission for writing to x. When a permission is used for a write access, it becomes an permission, and all foreign permissions become , canceling their reservations. Since z is adjacent to y, the write on line 5 is a foreign access.
@@
The second category of undefined behavior is an insufficient permission error. This type of error occurs when a reference or pointer with a read-only permission is used to write to memory. This occurs in Figure II-A on line 5, where the function ub writes through a pointer that was cast from an immutable reference ( &x) on line 2. Under Tree Borrows, this reference is given a permission, which is read-only. Casting a reference into a raw pointer does not change its permission, and neither does casting a pointer between const and mut.
The third type of undefined behavior that we observed is a framing error. We provide an example in Figure 1(c), where the function free receives a mutable reference. This reference encodes the promise that ownership will eventually be returned to the caller. However, this promise is broken when the reference is cast into a raw pointer and used to free the underlying allocation. In terms of Tree Borrows, when a reference-type value is passed as an argument to a function, it becomes protected, making it undefined behavior for it to transition to ---even if it is never accessed after this point. When a function returns, the permissions of its arguments are no longer protected.
@@ II-A1 Differences from Stacked Borrows
Tree Borrows replaces the Stacked Borrows [9] model, which uses a stack to track the active permissions to each location. Under Stacked borrows, a reference cannot be offset and used to access locations outside of the range it originally borrowed. Figure II-A shows an example of this access out-of-bounds error; the dereference on line 3 is invalid, since x only borrowed the first element of the tuple. This pattern is allowed under Tree Borrows; a valid permission can be used to access any location within the bounds of its allocation. Tree Borrows also delays asserting that the permission of a reference is valid until it is used for the first time, whereas Stacked Borrows asserts that a reference is valid immediately when it is created. However, Tree Borrows is less permissive than Stacked Borrows for certain aliasing patterns due to the semantics of its permission, justifying our second research question.
@@
Rust shares several other categories of undefined behavior with current C [16] and C++ [17] standards, such as accessing memory that has been freed, accessing beyond the bounds of an allocation, creating a data race, reading uninitialized memory, and accessing a value at an unaligned address [18]. However, C and C++ have different rules related to initialization, aliasing, and mutability that may conflict with Rust’s expectations.
@@ II-B1 Initialization
In both C and C++, accessing an uninitialized object produces an ‘‘indeterminate value’’, which is undefined behavior.222C23 §6.2.4.6-7; C++23 §6.7.4.1 However, there are a few exceptions. Static and thread-local variables are initialized immediately at the beginning of their lifetime, and both the unsigned char and byte types are allowed to store indeterminate values. It is also safe to create a reference to an uninitialized value.
Rust’s borrow checker prevents variables from being used until they are initialized, and no primitive types tolerate indeterminate values in safe contexts. However, Rust’s standard library provides the MaybeUninit<T> struct, which represents an instance of T that may or may not be initialized. Once an instance has been fully initialized, the unsafe function MaybeUninit<T>::assume_init() will unwrap the outer struct to produce the value inside. It is undefined behavior to call assume_init() if T is not fully initialized.
@@ II-B2 Aliasing & Mutability
Neither C nor C++ statically restricts aliasing, but the standards for each language allow implementations to make type-directed assumptions about aliasing. In both languages, it is considered undefined behavior for a pointer to one type to refer to a value of another type if the two types differ.333C23 §6.5.7; C++23 §7.2.1.11 However, differences do not include qualifiers, so two parameters of types int * and const int * can alias.
In Rust, both variables and references have distinct, connected capabilities for mutation. Variables are immutable by default, only mutable variables can be mutably borrowed, and only mutable references can be used for mutation. In C and C++, variables are mutable by default, and the capability to mutate memory is determined by the mutability of the object being pointed to, regardless of the pointer’s type.444C23 §6.7.4.1.7; C++23 §9.2.9.2.3 If a variable was declared as mutable, then a const pointer can be cast into a mutable pointer and used to write to it.
@@ II-B3 Provenance
Both C and C++ allow pointers to be converted to and from integers with a size equal to the word size of the current architecture.555C23 §6.3.2.3.5-6; C++23 §6.7.1.10.4-5 Implementations may track the provenance, or origin of pointers to inform compile-time optimizations [19], but neither of the current standards for C or C++ defines if or how provenance should be preserved across these conversions. Rust’s pointers have provenance [15], but the specifics of Rust’s provenance model are undecided.
Miri has multiple methods for handling provenance across pointer-to-integer conversion. By default, when a pointer is converted to an integer, its tag is added to a set of tags that have been ‘‘exposed’’ by pointer-to-integer conversion. When integers are converted back into pointers, their allocation identifier is reconstructed from their address, but they receive a wildcard provenance value. Under Stacked Borrows, when a memory access occurs using a wildcard tag, Miri eagerly interprets it as being equivalent to any tag in the stack that permits the access. The current implementation of Tree Borrows does not support this behavior yet; all accesses with wildcard tags are allowed and do not affect the state of the tree. Alternatively, developers can enable strict provenance, which treats integer-to-pointer conversion as an error.
@@ II-C Interoperation
Developers can access foreign functions and static variables using Rust’s foreign function interface. Declarations, or ‘‘bindings,’’ for each of these objects are written in
extern blocks, which take an optional qualifier to indicate the Application Binary Interface (ABI) used by the foreign library. Developers can also use the
extern keyword to change the ABI of a function defined in Rust so that it can be called from other languages. Only a subset of Rust’s types are guaranteed to be compatible with foreign ABIs. Both structs and enums must be annotated with the
#[repr(C)] attribute to ensure that their layout is compatible with C. The Rust compiler enforces the use of ABI-compatible types via lints, but these can be disabled. Bindings are not validated against their definitions, but tools such as bindgen [20] and CXX [21] can automatically generate bindings from headers.
Several dynamic analysis tools can find undefined behavior in multi-language Rust applications. Valgrind [22] is capable of detecting spatial and temporal memory errors, as well as incompatibilities with size and alignment. Many of the LLVM Project’s sanitizers are also compatible with Rust [23]. Each of these tools is available as a plugin for Cargo, Rust’s build tool. Miri can call foreign functions from natively compiled shared libraries using libffi [24], but it does not support passing certain argument types to native functions, and it cannot find undefined behavior that is triggered in foreign code.
@@ III Methodology
We seek to determine whether the differences between Rust and other languages lead to undefined behavior in practice. First, we analyzed all libraries published on crates.io---Rust’s central package repository---to find the subset with test cases that produce LLVM bitcode for C or C++ libraries during their build process. We describe this stage in Section III-A. Then, we created MiriLLI, which can run these test cases and detect undefined behavior. MiriLLI extends Miri to interoperate with LLI [25], an LLVM interpreter. Both Miri and LLI are included within the Rust toolchain; LLI is part of the LLVM backend. MiriLLI uses each interpreter to jointly execute programs defined across LLVM’s intermediate representation (LLVM IR) and Rust’s mid-level intermediate representation (MIR). In Section III-B, we outline the architecture of our tool and describe how we resolved the differences in semantics between each interpreter. In Section III-C, we describe our method for deduplicating test results to identify unique instances of undefined behavior and our approach to reporting bugs.